OpenStudy (anonymous):

Simplify the following polynomial and write the answer in standard form. (x^2+3x+2)+(2x^2-4x+1) a. -x+x^2+1 b. x^2-x+1 c. -x+3x^2+3 d. 3x^2-x+3

OpenStudy (the_fizicx99):

Again, except you're adding so just combine like terms. (x^2 + 2x^2) = ? (3x - 4x) = ? (2 + 1) = ?

OpenStudy (the_fizicx99):

Btw, x^2 is the same thing as saying 1x^2

OpenStudy (anonymous):

im thinking d?

OpenStudy (the_fizicx99):

Yeap, D!